[Brahms-l] Shift report 20050212 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat Feb 12 2005 - 16:05:33 EST
1. Came with nice beam, dumped at 12:00. Obtained new store at 14:20.

2. Runs:
FS@3A3450, MRS@40B2525
RunNo  trig2  trig3
13129 -  79k , 194k
13130 - 125k , 160k
13131 - 132k , 96k
FS@3B3450, MRS@40B2255
13136 - 108k , 60k 
move to
FS@4A2442, MRS@40B2255
RunNo -  Trig2   Trig3
13140 -   73k ,  58k
13144 -  171k ,  83k
13145 -   190k,  80k

Pedestals: 13141, 13142.

13146 - running.

3. Trig6 had some problem from last shift. Hiro came, and he did some test on it. Seems everything came back.



_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Feb 12 16:05:55 2005

This archive was generated by hypermail 2.1.8 : Sat Feb 12 2005 - 16:06:03 EST